A force produces an acceleration of 5m/s2 in a body and same forces produces acceleration of 15m/s2 in another body.the acceleration produced by the same fore applied to combination of system of two bodies will be

Grade:12

1 Answers

Arun
6 years ago
Dear Aniket
 
M = m1 + m2
F/a = F/a1 + F/a2
 
1/a = 1/a1 + 1/a2
a = a1 * a2/(a1 + a2)
a = 5*15/20
a = 75/20 = 3.75 m/sec²